### v1

@@ 5.2 Tensor $\ell_{2,p}$ norm for group sparsity
The matrix $\ell_{2,p}$ norm is a nonconvex and nonsmooth function. And it has been applied to image processing [20], machine learning [23, 11], feature selection [19, 31], multi-view classification [33], etc. To measure the linear structural sparsity of the sparse noise tensor $\mathcal{S}$, we extend the matrix $\ell_{2,p}$ norm for group sparsity to its tensor form. As the stripes and dead lines often align the first dimension, we define the tensor $\ell_{2,p}$ $(0<p<1)$ norm in the form of (2), which is exactly the matrix $\ell_{2,p}$ norm of the unfolding matrix along the first dimension.
@@
$s_{:i_{2}i_{3}}^{k+1}=\Gamma_{\mu}\left(\left\|\sqrt{w_{:i_{2}i_{3}}}\odot\tilde{s}^{k}_{:i_{2}i_{3}}\right\|_{2}\right)\tilde{s}^{k}_{:i_{2}i_{3}},$
where $\tilde{\mathcal{S}}^{k}=\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D}\right)$, and $\tilde{s}^{k}_{:i_{2}i_{3}}$ and $\sqrt{w_{:i_{2}i_{3}}}$ are the $(i_{2},i_{3})$-th mode-$1$ fibers of $\tilde{\mathcal{S}}^{k}$ and $\sqrt{\mathcal{W}_{\mathbf{R}}}$, respectively.
@@ 6.1 Simulated data experiments
(a) Washington DC Mall
(b) Xiong-An
Figure 3: Original HSIs of two datasets. (a) Part of Washington DC Mall dataset (R:17, G:36, B:46); (b) Xiong-An dataset (R:71, G:110, B:120).
To simulate the noisy HSI data, Gaussian noise, stripes, or dead lines are added to the normalized clean HSI data under the following cases:
  • Case 1: Gaussian noise with a mean of zero and a standard deviation of $0.1$ is added to all the bands. And then all the bands are selected and stripes with a density of $30\%$ and a standard deviation of $0.2$ are added to each band.
  • Case 2: Gaussian noise with a mean of zero and a standard deviation of $0.1$ is added to all the bands. And then $11-40$, $71-100$, $121-128$ bands are selected and stripes with a density of $20\%$ and a standard deviation of $0.2$ are added to each band.
  • Case 3: Gaussian noise with a mean of zero and a standard deviation of $0.2$ is added to all the bands. And then $25\%$ of the bands are randomly selected and dead lines with a density of $5\%$ are added to each band.
@@
The numerical results of simulated data experiments for case 1, case 2, and case 3 are presented in Tables 1 and 2 for Washington DC Mall and Xiong-An datasets, respectively. Numerical results in bold font indicate the best performance of the indicator in the current case. It can be observed from Tables 1 and 2 that the proposed NLTL2p method outperforms other methods almost in terms of all the evaluation metrics. For example, in case 1 of Xiong-An dataset, the MPSNR value of the HSI restored by the NLTL2p method is 1.84 dB larger than the MPSNR value of the second best method, that is, the LRTFL0 method.
Table 1: Numerical results tested on Washington DC Mall dataset
@@
1
MPSNR
14.90 30.49 30.29 25.92 26.47 16.35 25.27
MSSIM
0.324 0.928 0.923 0.795 0.817 0.399 0.794
MFSIM
0.653 0.949 0.957 0.902 0.907 0.724 0.885
ERGAS
518.02 92.41 92.45 151.04 138.47 438.24 159.34
2
MPSNR
16.72 32.14 30.46 31.45 27.82 22.72 26.97
MSSIM
0.428 0.939 0.920 0.912 0.854 0.639 0.844
MFSIM
0.703 0.962 0.933 0.950 0.921 0.817 0.916
ERGAS
419.52 91.63 78.65 83.10 105.07 330.58 119.73
3
MPSNR
13.90 30.66 29.83 28.01 25.44 20.80 25.62
MSSIM
0.390 0.943 0.924 0.895 0.801 0.762 0.842
MFSIM
0.680 0.963 0.929 0.943 0.897 0.879 0.913
ERGAS
610.91 90.35 93.19 141.32 148.89 449.91 147.69
Table 2: Numerical results tested on Xiong-An dataset
Case
Index
Noisy
NLTL2p
LRTFL0
SNLRSF
LRTD
BM4D
QRNN3D
1
MPSNR
14.49 33.57 31.73 26.26 30.02 15.88 26.53
MSSIM
0.102 0.864 0.810 0.562 0.710 0.151 0.633
MFSIM
0.482 0.928 0.915 0.829 0.883 0.575 0.852
ERGAS
382.96 43.96 55.28 112.07 66.10 328.21 97.50
2
MPSNR
17.19 33.43 33.21 30.25 30.71 24.29 28.27
MSSIM
0.159 0.875 0.837 0.719 0.730 0.468 0.694
MFSIM
0.562 0.928 0.924 0.885 0.894 0.722 0.880
ERGAS
306.56 45.12 48.82 80.07 60.80 234.71 82.38
3
MPSNR
13.90 31.48 30.89 28.39 27.96 21.30 26.13
MSSIM
0.238 0.912 0.858 0.828 0.765 0.692 0.755
MFSIM
0.634 0.954 0.897 0.924 0.881 0.831 0.900
ERGAS
435.95 58.47 56.64 103.41 83.80 321.25 101.38
For visual quality comparison, the HSIs restored by different methods are presented in Fig. 4-7, with a subregion marked by a white box and enlarged in a white box. The NLTL2p method achieves state-of-the-art performance for HSI denoising and destriping, while the competing methods seem to fail to remove noises or to restore HSIs with high quality. For example, for removing mixed noise, the BM4D method fails to remove stripes even if the HSI only contains few stripes as shown in Fig. 5(5(g)); the SNLRSF and QRNN3D methods fail to remove the stripes and dead lines, when the noise level is high, for example, in Fig. 4(4(e))(4(h)) and Fig. 6(6(e))(7(h)). Also, regarding the quality of the restoration, the HSIs restored by the LRTD method look a little noisy and blurry, for example, in Fig. 4(4(f)); the HSIs restored by the LRTFL0 method looks oversmooth with some details missing, for example, in Fig. 6(6(d)) and Fig. 7(7(d)).
@@
Figure 4: Comparison of HSIs (R:3, G:43, B:75) restored by different methods from Washington DC Mall in case 1. The PSNR value for each restored HSI: (a) Noisy image (14.90 dB); (c) NLTL2p (ours) (30.49 dB); (d) LRTFL0 (30.29 dB); (e) SNLRSF (25.92 dB); (f) LRTD (26.47 dB); (g) BM4D (16.35 dB); (h) QRNN3D (25.27 dB).
@@
Figure 5: Comparison of HSIs (R:16, G:70, B:100) restored by different methods from Washington DC Mall in case 3. The PSNR value for each restored HSI: (a) Noisy image (13.90 dB); (c) NLTL2p (ours) (30.66 dB); (d) LRTFL0 (29.83 dB); (e) SNLRSF (28.01 dB); (f) LRTD (25.44 dB); (g) BM4D (20.80 dB); (h) QRNN3D (25.62 dB).
@@
Figure 6: Comparison of HSIs (R:30, G:55, B:115) restored by different methods from Xiong-An in case 1. The PSNR value for each restored HSI: (a) Noisy image (14.49 dB); (c) NLTL2p (ours) (33.57 dB); (d) LRTFL0 (31.73 dB); (e) SNLRSF (26.26 dB); (f) LRTD (30.02 dB); (g) BM4D (15.88 dB); (h) QRNN3D (26.53 dB).
@@
Figure 7: Comparison of HSIs (R:38, G:87, B:120) restored by different methods from Xiong-An in case 3. The PSNR value for each restored HSI: (a) Noisy image (13.90 dB); (c) NLTL2p (ours) (31.48 dB); (d) LRTFL0 (30.89 dB); (e) SNLRSF (28.39 dB); (f) LRTD (27.96 dB); (g) BM4D (21.30 dB); (h) QRNN3D (26.13 dB).
@@ 6.2 Real data experiments
In this subsection, we test the proposed method and the competing methods on two real HSI datasets containing mixed noise. The test images are subimages of size $128\times 128\times 128$ randomly obtained from the HYDICE Urban333 http://www.erdc.usace.army.mil/Media/Fact-Sheets/Fact-Sheet-Article-View/Article/610433/hypercube/ $(307\times 307\times 210)$ and EO-1 Hyperion [^4] http://www.lmars.whu.edu.cn/prof_web/zhanghongyan/resource/noise_EOI.zip $(400\times 200\times 166)$, which are shown in Fig. 8. A selected band of the HSI restored by each method is presented in Fig. 9 and Fig. 10 for HYDICE Urban dataset and EO-1 Hyperion dataset, respectively. It can be observed that the proposed NLTL2p method can remove the stripes while preserving the image details. However, the LRTFL0, LRTD, BM4D and QRNN3D methods are unable to eliminate the stripes when the band is contaminated by heavy mixed noise as shown in Fig. 9; and the SNLRSF, LRTD, BM4D and QRNN3D methods remove not only the noise but also some structural details of the HSI as shown in Fig. 10.
(a) HYDICE Urban
(b) EO-1 Hyperion
Figure 8: Real HSIs of two datasets. (a) HYDICE Urban dataset (R:61, G:98, B:170); (b) EO-1 Hyperion dataset (R:101, G:114, B:160).
@@
6.3 Parameter analysis
For the proposed NLTL2p method, we set block matching with $(m_{1},m_{2},m_{3})=(25,128,128)$ and $(n_{1},n_{2},n_{3})=(25,2,32)$; we set model parameters $\delta\in[0.4,0.6]$, $w_{j}=10^{2}$, $p=0.1$, and $\gamma\in[100,200]$; and we set algorithm parameters $\alpha_{\mathcal{S}}=\alpha_{X}=\alpha_{\mathcal{G}}=0.001$. In practice, to accelerate the algorithm, we update $\mathbf{R}$ via block matching on FBBs for several iterations in the beginning and then fix it for the remaining iterations. Also, we perform the updates of $[X_{i}]$ and $[\mathcal{G}]$ three times as inner iterations for boosting the nonlocal low-rank regularization process. Note that these practical strategies will not affect the convergence of the algorithm. And Algorithm 1 stops if $\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}/\|\mathcal{L}^{k}\|_{F}\leq 0.01$.
To further analyze the parameters in the proposed model, we use case 3 of Washington DC Mall for testing. First, we test the proposed NLTL2p method with different values of $\delta$ and $\gamma$. The plot of MPSNR values vs parameters is presented in Fig. 11(11(a)). When the value of $\gamma$ is fixed, the change of MPSNR over $\delta$ is not obvious. When the value of $\delta$ is fixed, the MPSNR value is larger than 29 dB if $\gamma\in[100,200]$.
Second, we test the proposed NLTL2p method with different values of $p$ and $\gamma$ and plot the MPSNR values vs parameters in Fig. 11(11(b)). When the value of $\gamma$ is fixed, the MPSNR value is larger if $p$ is smaller. When the value of $p$ is less than $0.5$, the best MPSNR value is achieved if $\gamma\in[100,200]$.
In summary, the proposed method is not very sensitive to $\delta$ and can achieve great performance when $\gamma$ is between $100$ and $200$, and $p$ is close to 00.
(a) Plot on $\delta$ and $\gamma$.
(b) Plot on $p$ and $\gamma$
Figure 11: Plots of the MPSNR values obtained by our proposed method vs several parameters.

### v2

@@ 5.3 Tensor $\ell_{2,p}$ norm for group sparsity
To remove stripes in HSIs, we measure the linear structural sparsity of the sparse noise tensor $\mathcal{S}$ using the tensor $\ell_{2,p}$ norm. The tensor $\ell_{2,p}$ norm for group sparsity is extended from the matrix $\ell_{2,p}$ norm, which has been applied to image processing [24], machine learning [27, 15], feature selection [23, 35], multi-view classification [37], etc. As the stripes and dead lines often align the first dimension, we define the tensor $\ell_{2,p}$ $(0<p<1)$ norm in the form of (2), which is exactly the matrix $\ell_{2,p}$ norm of the unfolding matrix along the first dimension.
@@
$s_{:i_{2}i_{3}}^{k+1}=\Gamma_{\tilde{\gamma}}\left(\left\|\tilde{s}^{k}_{:i_{2}i_{3}}\right\|_{2}\right)\tilde{s}^{k}_{:i_{2}i_{3}},$
where $\tilde{\mathcal{S}}^{k}=\mathcal{S}^{k}-\tilde{\alpha}_{\mathcal{S}}\left(\mathcal{S}^{k}+\mathcal{L}^{k}-\mathcal{D}\right)$, and $\tilde{s}^{k}_{:i_{2}i_{3}}$ is the $(i_{2},i_{3})$-th mode-$1$ fibers of $\tilde{\mathcal{S}}^{k}$.
5.4 Two-phase MLTL2p method
We extend model (1) to a multi-scale low-rank model and apply the P-BCD method in a two-phase fashion for HSI denoising and destriping.
The spectral correlation and spatial nonlocal self-similarity in HSIs imply multi-scale low-rankness, which includes global low-rankness, local block low-rankness, and nonlocal self-similar tensor low-rankness. In the following, we propose a multi-scale low-rank model that imposes low-rankness on global, local, and nonlocal tensors extracted via $\mathbf{R}_{g}$, $\mathbf{R}_{l}$, and $\mathbf{R}_{nl}$, respectively.
(31) $\begin{split}\min_{\begin{subarray}{c}[{X}_{i}]_{g},[{X}_{i}]_{l},[{X}_{i}]_{nl},\\ [\mathcal{G}]_{g},[\mathcal{G}]_{l},[\mathcal{G}]_{nl},\mathcal{S},\mathcal{L}\end{subarray}}\quad&\frac{1}{2}\|\mathcal{L}+\mathcal{S}-\mathcal{D}\|_{F}^{2}+\gamma\|\mathcal{S}\|^{p}_{2,p}\\ &+\|[\mathcal{G}]_{g}\|_{1,w}+\frac{\delta_{g}}{2}\|\mathbf{R}_{g}(\mathcal{L})-[\mathcal{G}]_{g}\times_{1}[X_{1}]_{g}\times_{2}[X_{2}]_{g}\times_{3}[X_{3}]_{g}\|^{2}_{F}\\ &+\|[\mathcal{G}]_{l}\|_{1,w}+\frac{\delta_{l}}{2}\|\mathbf{R}_{l}(\mathcal{L})-[\mathcal{G}]_{l}\times_{1}[X_{1}]_{l}\times_{2}[X_{2}]_{l}\times_{3}[X_{3}]_{l}\|^{2}_{F}\\ &+\|[\mathcal{G}]_{nl}\|_{1,w}+\frac{\delta_{nl}}{2}\|\mathbf{R}_{nl}(\mathcal{L})-[\mathcal{G}]_{nl}\times_{1}[X_{1}]_{nl}\times_{2}[X_{2}]_{nl}\times_{3}[X_{3}]_{nl}\|^{2}_{F}\\ \mbox{s.t. }&[{X}_{i}]_{g}^{\top}[{X}_{i}]_{g}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\\ &[{X}_{i}]_{l}^{\top}[{X}_{i}]_{l}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3,\\ &[{X}_{i}]_{nl}^{\top}[{X}_{i}]_{nl}=[\mathbf{I}_{n_{i}}],\quad i=1,2,3.\end{split}$
And the P-BCD method for solving model (31) is presented in Algorithm 2.
Algorithm 2 P-BCD algorithm for model (31)
1:  Initialize $(\mathcal{S}^{0},[{X}_{i}^{0}]_{g},[{X}_{i}^{0}]_{l},[{X}_{i}^{0}]_{nl},[\mathcal{G}^{0}]_{g},[\mathcal{G}^{0}]_{l},[\mathcal{G}^{0}]_{nl},\mathcal{L}^{0})$ with $[{X}_{i}^{0}]_{g},[{X}_{i}^{0}]_{l},[{X}_{i}^{0}]_{nl}$ being independently orthogonal;
2:  Set the tensor extraction operators $\mathbf{R}_{g}$, $\mathbf{R}_{l}$, and $\mathbf{R}_{nl}$;
4:  Set $k=0$.
5:  repeat
6:     Compute $\mathcal{S}^{k+1}$ by (7);
7:     Compute $[X_{i}^{k+1}]_{g},[{X}_{i}^{k+1}]_{l},[{X}_{i}^{k+1}]_{nl}$ by (9), $i=1,2,3$;
8:     Compute $[\mathcal{G}^{k+1}]_{g},[\mathcal{G}^{k+1}]_{l},[\mathcal{G}^{k+1}]_{nl}$ by (11);
9:     Compute $\mathcal{L}^{k+1}$ by
$\mathcal{L}^{k+1}=$ $\left(\delta_{g}\mathcal{W}_{\mathbf{R}_{g}}+\delta_{l}\mathcal{W}_{\mathbf{R}_{l}}+\delta_{nl}\mathcal{W}_{\mathbf{R}_{nl}}+\operatorname{Id}\right)^{-1}$
$\odot\Big{(}\delta_{g}\mathbf{R}_{g}^{\top}([\mathcal{Y}^{k+1}]_{g})+\delta_{l}\mathbf{R}_{l}^{\top}([\mathcal{Y}^{k+1}]_{l})$
$+\delta_{nl}\mathbf{R}_{nl}^{\top}([\mathcal{Y}^{k+1}]_{nl})+(\mathcal{D}-\mathcal{S}^{k+1})\Big{)};$
10:     $k\leftarrow k+1$.
11:  until the stopping criterion is met.
11:  $(\mathcal{S}^{k},[{X}_{i}^{k}]_{g},[{X}_{i}^{l}]_{l},[{X}_{i}^{k}]_{nl},[\mathcal{G}^{k}]_{g},[\mathcal{G}^{k}]_{l},[\mathcal{G}^{k}]_{nl},\mathcal{L}^{k})$
To enhance the efficiency and effectiveness of the MLTL2p method, we implement a two-phase scheme on the multi-scale low-rank model and the P-BCD method. In the first phase, we employ only the global and local low-rank regularization, which is sufficient for identifying stripes and dead lines. In particular, we use Algorithm 2 to solve model (31) with the parameter $\delta_{nl}$ set to zero. In the second phase, we utilize the nonlocal low-rank regularization along with the global and local low-rank regularization. As the estimated HSI from the first phase has already removed most of the noise, we can perform only one block matching to obtain a fixed nonlocal block extraction $\mathbf{R}_{nl}$, which can significantly reduce the computational cost. By applying Algorithm 2 to solve model (31) in this phase, the algorithm can refine additional details in the HSI, thereby improving the quality of the restored HSI.
@@ 6.1 Simulated data experiments
To simulate the noisy HSI data, Gaussian noise, stripes, or dead lines are added to the normalized clean HSI data using the procedure in [9] under the following cases:
  • Case 1: Gaussian noise with a mean of zero and a standard deviation of $0.1$ is added to all the bands. Additionally, bands $45-60$ and $105-120$ (i.e., $25\%$ of the bands) are selected to impose stripe noise. Stripe noise with an intensity (mean absolute value of pixels) of $0.2$ and a percentage of $10\%$ per band (i.e., $2.5\%$ of the HSI) is added to the selected bands.
  • Case 2: Gaussian noise with a mean of zero and a standard deviation of $0.1$ is added to all the bands. Additionally, all the bands are selected to impose stripe noise. Stripe noise with an intensity of $0.2$ and a percentage of $10\%$ per band (i.e., $10\%$ of the HSI) is added to all the bands.
  • Case 3: Gaussian noise with a mean of zero and a standard deviation of $0.1$ is added to all the bands. Additionally, all the bands are selected to impose dead lines. Dead lines with a percentage of $5\%$ per band (i.e., $5\%$ of the HSI) are added to the selected bands.
  • Case 4: Gaussian noise with a mean of zero and a standard deviation randomly varying from $0.1$ to $0.2$ is added to all the bands. Additionally, $25\%$ of the bands are selected from band $1-64$ to impose stripe noise with an intensity of $0.2$ and a percentage of $10\%$ per band; $12.5\%$ of the bands are selected from band $65-128$ to impose dead lines with a percentage of $5\%$ per band.
@@
For the proposed MLTL2p method, we set algorithm parameters $\alpha_{\mathcal{S}}=0.1$, and $\alpha_{X}=\alpha_{\mathcal{G}}=0.01$ for Algorithm 2. In the first phase, we set model parameters $w_{j}=10^{2}$, $p=0.1$, and $\delta_{nl}=0$; we set $\gamma=0.8$ for case 1 and 2 and $\gamma=1$ for case 3 and 4; for $\mathbf{R}_{g}$, we set $\delta_{g}=1$, block size $[128,128,128]$ and block rank $[102,102,3]$; for $\mathbf{R}_{l}$, we set $\delta_{l}=1$, block size $[32,32,32]$ and block rank $[26,26,2]$. Algorithm 2 stops at iteration $10$. In the second phase, we set model parameters $w_{j}=10^{2}$, $p=0.1$; we set $\gamma=1.76$ for case 1 and 2 and $\gamma=2.2$ for case 3 and 4; for $\mathbf{R}_{g}$, we set $\delta_{g}=3$, block size $[128,128,128]$ and block rank $[102,102,5]$; for $\mathbf{R}_{l}$, we set $\delta_{l}=3$, block size $[32,32,32]$ and block rank $[26,26,3]$; for $\mathbf{R}_{nl}$, we set $\delta_{nl}=60/\operatorname{median}(\mathcal{W}_{\mathbf{R}})$, block size $[36,128,128]$ and block rank $[32,43,5]$. Algorithm 2 stops if $\|\mathcal{L}^{k+1}-\mathcal{L}^{k}\|_{F}/\|\mathcal{L}^{k+1}\|_{F}\leq 0.005$ and $\|\mathcal{S}^{k+1}-\mathcal{S}^{k}\|_{F}/\|\mathcal{S}^{k+1}\|_{F}\leq 0.005$. For the competing methods, the parameters are manually tuned according to the settings provided in their articles to ensure optimal performance.
Table 1: Average numerical results tested on simulated data
@@
1 MPSNR 19.63 39.35 37.02 39.10 30.97 33.04 31.56
MSSIM 0.276 0.969 0.940 0.965 0.807 0.858 0.856
MFSIM 0.613 0.981 0.968 0.978 0.907 0.918 0.926
ERGAS 0.426 0.041 0.057 0.043 0.114 0.117 0.110
2 MPSNR 18.52 39.03 35.99 34.97 30.64 26.00 31.05
MSSIM 0.245 0.967 0.929 0.915 0.797 0.650 0.840
MFSIM 0.580 0.980 0.962 0.953 0.902 0.826 0.918
ERGAS 0.470 0.043 0.064 0.072 0.120 0.214 0.118
3 MPSNR 14.46 35.51 34.91 26.24 29.32 15.82 30.88
MSSIM 0.234 0.929 0.920 0.704 0.782 0.576 0.829
MFSIM 0.572 0.967 0.957 0.856 0.903 0.764 0.916
ERGAS 0.600 0.074 0.086 0.170 0.133 0.540 0.119
4 MPSNR 16.10 36.49 34.64 35.33 26.47 23.99 30.35
MSSIM 0.176 0.945 0.909 0.925 0.579 0.503 0.819
MFSIM 0.513 0.966 0.949 0.955 0.804 0.759 0.906
ERGAS 0.593 0.061 0.099 0.078 0.208 0.325 0.125
Mtime -- 299.0 265.9 302.2 207.7 104.9 0.6
The average numerical results for each case over all the simulated HSIs are presented in Table 1. The results in bold font indicate the best performance of the evaluation metric in the current case. It can be observed that the proposed MLTL2p method outperforms other methods in terms of all the evaluation metrics. For example, in case 2, the MPSNR value of the HSI restored by the MLTL2p method is 3 dB larger than the MPSNR value of the second best method, that is, the LRTFL0 method.
For visual quality comparison, the HSIs restored by different methods are presented in Fig. 5-8, with a subregion marked by a white box and enlarged in a white box. The MLTL2p method achieves state-of-the-art performance for HSI denoising and destriping, while the competing methods seem to fail to remove noises or restore HSIs with high quality. For example, for removing mixed noise, the BM4D method fails to remove stripes even if the HSI only contains few stripes as shown in Fig. 5(6(g)); the SNLRSF and QRNN3D methods may not remove the stripes and dead lines, when the noise percentage is high, for example, in Fig. 7(5(e))(5(h)). Also, regarding the quality of the restoration, the HSIs restored by the LRTFL0 method may contain staircase artifacts, e.g., in Fig. 5(5(d)); the HSIs restored by the LRTD method contain blur and artifacts, e.g., in Fig. 5(5(f)); the HSIs restored by the QRNN3D method look noisy, e.g., in Fig. 5(5(h)).
@@
Figure 5: Comparison of HSIs (R:112, G:70, B:20) restored by different methods from Chikusei in case 1. The PSNR value for each restored HSI: (a) Noisy image (19.63 dB); (c) MLTL2p (ours) (39.57 dB); (d) LRTFL0 (37.05 dB); (e) SNLRSF (38.70 dB); (f) LRTD (31.82 dB); (g) BM4D (33.79 dB); (h) QRNN3D (32.20 dB).
@@
Figure 6: Comparison of HSIs (R:118, G:78, B:38) restored by different methods from Xiong-An in case 2. The PSNR value for each restored HSI: (a) Noisy image (18.52 dB); (c) MLTL2p (ours) (38.82 dB); (d) LRTFL0 (35.74 dB); (e) SNLRSF (34.76 dB); (f) LRTD (31.90 dB); (g) BM4D (26.16 dB); (h) QRNN3D (31.01 dB).
@@
Figure 7: Comparison of HSIs (R:60, G:27, B:17) restored by different methods from Washington DC Mall in case 3. The PSNR value for each restored HSI: (a) Noisy image (14.66 dB); (c) MLTL2p (ours) (35.28 dB); (d) LRTFL0 (34.62 dB); (e) SNLRSF (26.65 dB); (f) LRTD (28.52 dB); (g) BM4D (16.15 dB); (h) QRNN3D (30.01 dB).
@@
Figure 8: Comparison of HSIs (R:107, G:73, B:21) restored by different methods from Xiong-An in case 4. The PSNR value for each restored HSI: (a) Noisy image (16.02 dB); (c) MLTL2p (ours) (36.66 dB); (d) LRTFL0 (34.95 dB); (e) SNLRSF (35.46 dB); (f) LRTD (25.91 dB); (g) BM4D (23.12 dB); (h) QRNN3D (29.54 dB).
@@ 6.2 Real data experiments
In this subsection, we test the proposed method and the competing methods on two real HSI datasets containing mixed noise. The test images are subimages of size $128\times 128\times 128$ randomly obtained from the HYDICE Urban444 http://www.erdc.usace.army.mil/Media/Fact-Sheets/Fact-Sheet-Article-View/Article/610433/hypercube/ $(307\times 307\times 210)$ and EO-1 Hyperion [^5] http://www.lmars.whu.edu.cn/prof_web/zhanghongyan/resource/noise_EOI.zip $(400\times 200\times 166)$. A selected band of the HSI restored by each method is presented in Fig. 9 and Fig. 10 for HYDICE Urban dataset and EO-1 Hyperion dataset, respectively. It can be observed that the proposed MLTL2p method can remove the stripes while preserving the image details. However, the LRTFL0, LRTD, BM4D and QRNN3D methods are unable to eliminate the stripes when the band is contaminated by heavy mixed noise as shown in Fig. 9; and the SNLRSF, LRTD, BM4D and QRNN3D methods remove not only the noise but also some structural details of the HSI as shown in Fig. 10.